Given the following measurements from John’s health assessment, answer the question given:

            End diastolic volume = 105 ml

            End systolic volume = 30 ml

            Heart rate = 70 bpm

What is John’s maximum heart rate? (hint: subtract his age from 220).

(0.1pts)

Question 82 - Given the following measurements from John’s health assessment, answer the question given:             End diastolic volume = 105 ml             End systolic volume = 30 ml             Heart rate = 70 bpm What is John’s maximum heart rate? (hint: subtract his age from 220).

Select 115 bpm as your answer

115 bpm

Select 220 bpm as your answer

220 bpm

Select 200 bpm as your answer

200 bpm

Select 150 bpm as your answer

150 bpm
Answer #1

To calculate John"s maximum heart rate , we have to subtract his age from 220.

The age is not mentioned in the question,but we can find the answer by eliminating the options.

A-- 115 bpm

So age has to be = 220-115 = 105 years

Not possible.

B-- 220 bpm

Age = 0 years

Not possible

C-- 200 bpm

Age = 220-200=20 years

D-- 150 bpm

Age = 220-150= 70 years

Seeing the stroke volume of 75 ml (105-30=75 ml), option C (20 years ) is more likely.

So answer is C.